OpenStudy (anonymous):

Solve the equation by the square root property. (x + 1)2 = - 6

OpenStudy (anonymous):

\[ \sqrt{(x+1)^2} = \sqrt{-6} \\ |x+1| = \sqrt{6}\cdot\sqrt{-1}\\ \pm(x+1) = \sqrt{6}\cdot\sqrt{-1} \]

OpenStudy (anonymous):

i don't understand can you please explain it??